Weekly Update

10-Apr-2018

It’s been below zero for the past few days so we haven’t been able to replace the frames that made way for the package off bees.

Today the frames were replaced and the comb that the bees formed in the empty space was broken out. Eggs were found in each of the hives.

Bees are still eating honey, next week bee feed will be added to the hives. There are still some days in the next week that will be below zero.

April 10th 2018
<p>5-Apr-2018</p>

<p>Today our bees were picked up from Charlie Newton.  The bees from the 3lb. were introduced to the hives and one entrance left open (only one bee in/out at a time) to allow the stragglers from the package to enter the hive.</p>

<p>The temperature is low, nightly lows are in the teens.  Each hive was provided with, a full frame of honey and a full pollen pattie.  12 gallons of bee feed was made and is on hand to start supplementing tomorrow.</p>

New stands were made that stand a foot above the ground and anchored in the soil about a foot.

5-Apr-2018

Today our bees were picked up from Charlie Newton. The bees from the 3lb. were introduced to the hives and one entrance left open (only one bee in/out at a time) to allow the stragglers from the package to enter the hive.

The temperature is low, nightly lows are in the teens. Each hive was provided with, a full frame of honey and a full pollen pattie. 12 gallons of bee feed was made and is on hand to start supplementing tomorrow.

New stands were made that stand a foot above the ground and anchored in the soil about a foot.
April 5th 2018

Bee Class

19-Jan-2018

I am teaching a basic backyard beekeeping class in April.  The course covers Bee Biology, Hive maintenance, and Pests.  The main source for the material is, “The Hive and the Honeybee” 2015 ed.

January 20th 2018

Expansion

19-Jan-2018

Our little bee yard is expanding to 3 hives, the previous Medu (Slovene) hive is being resurrected and two new hives are being added, Mel (Latin) and Medlin (Elvish).  All hives are Apimaye Thermo Langstroth Hives.  Three 3lb packages have been ordered from Newton’s Own Hive & Honeybee Products.  The colonies are planned to start in April.

January 20th 2018

Medu Update

13-Nov-2017

Medu Bees were no longer taking the 2:1 syrup because it was below 50° for the past couple of weeks. We switched to candy instead and while switching food sources we noticed the colony only had a few thousand bees left. We checked the hive yesterday and found that all of the bees were dead and they looked like they died while performing there tasks, helping new bees out of the cells, feeding, etc… Many bees were found in the hive entrance and most in the bottom board. It looks like dysentery, but not nosema.

The hive was dismantled and the comb was cut out and the frames and body were cleaned. We expect that it was the changing food source that caused the dysentery, but we are not positive what killed the colony.

November 13th 2017

Bees Preparing for Fall

The undertaker Bees are working full tilt cleaning out dead and dying Bees. There are also drones being surrounding the entrance.

September 14th 2017

Social “Bees”

We attended a talk given by the Iowa Honey Queen and networked with local beekeepers.



We plan on attending the annual Iowa Honey Producers Association 2-day conference, to network and further learn the tricks of the trade.

September 2nd 2017

Wintering Preparation Started

Over the past couple of weeks we extracted 5 frames and removed empty frames so we could remove Medu-Super 3 from the hive. The very dry summer had taken its toll on flower production and so we aren’t expecting any late up swing in honey.

September 2nd 2017
<p>In preparation of adding another hive next year, we named and labeled the current hive, Medu.  Medu is Slovene for honey.  </p>

<p>The first honey we harvested last month was two frames out of Super 3.  On Friday we extracted two more frames out of Super 3.  Total we harvested ~1.5 gallons.</p>

In preparation of adding another hive next year, we named and labeled the current hive, Medu. Medu is Slovene for honey.

The first honey we harvested last month was two frames out of Super 3. On Friday we extracted two more frames out of Super 3. Total we harvested ~1.5 gallons.

August 6th 2017
<p>This is the first bottle of honey harvested from our first hive.</p>

This is the first bottle of honey harvested from our first hive.

July 19th 2017
<p>There has been a constant flow of bees at the fountain for the past month.  They have also been consuming ~6 liters of bee feed a week from the top feeder on the hive.</p>

There has been a constant flow of bees at the fountain for the past month. They have also been consuming ~6 liters of bee feed a week from the top feeder on the hive.

July 12th 2017
tumblr photo tumblr photo
<p>Full hive boxes.  The bees are bearding a little at the entrance.  At night the bearding covers the landing platform and the entrance with hanging bees.</p>

Full hive boxes. The bees are bearding a little at the entrance. At night the bearding covers the landing platform and the entrance with hanging bees.

July 11th 2017

Update

10-Jul-2017

The bees have been bearding at the hive entrance for 3 days. It has been in the mid to high 90’s for the past couple of weeks. I added two more supers and split the brood boxes frames. The original super was mostly combed and now the two brood boxes and the two upper supers have five fully combed frames with honey and five empty frames with foundations. The queen excluder was removed in the middle of June.

July 11th 2017

Weekly Update

02-Jun-2017

The first super was placed on top of the hive and sperated by a queen excluder. The bee feed was refilled.

June 6th 2017

Weekly Update

20-May-2017

It’s been raining for the past couple of weeks. We inspected the hive for infestation and found hundreds of ants in the top feeder and in the bottom tray. We also found small hive beetles and their larva in the hive. We created a new base with a wide and completely closed bottom. The floor of the new base was smeared with Tanglefoot.

While in the hive we noticed that about 50% of the second brood box was drawn with comb.

May 20th 2017

Weekly Update

29-Apr-2017

It’s been raining for over a week and likely continue to rain for the next week. I refilled the liquid bee feed (it seems that a few bees ended up getting stuck and dying in the feed). I added a pollen pattie between the second brood box and top feeder to encourage comb production in the second brood box.

April 29th 2017
<p>Newly opened hive.</p>

Newly opened hive.

April 23rd 2017

Weekly Update

21-Apr-2017
We checked on comb production in the second brood box. There was 2 frames that had a small amount of comb.
We also surrounded the hive with Tangle Foot to prevent carpenter ants from entering the hive.

April 22nd 2017

Hive Expansion

17-Apr-2017

The frames were 65% drawn. We put another brood box on top full of ten frames with 100% Beeswax foundations.

April 17th 2017

Weekly Update

14-Apr-2017

We opened the hive to check on comb production. Five of the frames where completely or nearly completely drawn out with comb. One of those frames had honey and brood.
The entrance was opened to 100% including drone entrance. The pollen pattie was ~60% consumed and the liquid bee feed was refilled with a Honey be Healthy (HBH) formula;

1:1 sugar to water @ 2 kg each
2 ml of spearmint oil
2 ml of lemongrass oil
0.5 g of lecithin powder

April 14th 2017

Blue Bottle Flies

08-Apr-2017

Blue Bottle Flies were on and surrounding the hive in mass.  The under tray was checked and the hive was inspected for American Foul Brood (AFB) and European Foul Brood (EFB).  No negative signs were found of hive collapse.  While inspecting the hive, there was comb production starting.

April 9th 2017

Opening Day

07-Apr-2017

The queen was released by the bees and we removed her cage.  The hive entrance was opened to ~25% open.  The bees found a flowering bush immediately and filled their pollen baskets.  

April 9th 2017

Start of Hive

06-Apr-2017

We introduced a three pound package into an Apimaye langstroth hive.  The queen was placed in the hive with a marshmallow plug.  The top feeder was filled with liquid feed and a 15% pollen patty was placed on top of the frames.  The entrance was closed.  

April 9th 2017